5. Compare the following two alternatives by the IRR method, given MARR of 8%/year. Is the incement in cost form A to B justiPlease don't provide the answer by excel calculation. I need to understand the formula and step by step.Thank you.
Answer #1

Here we are required to calculate the IRR of the two alternatives.

First of all you are required to write the equation for NPW of the alternatives. Alternative A NPW can be written as

\large NPW_A = - 510,000 + 145,000(P/A,i\%,11) + 10,000(P/F,i\%,11)

\large \implies NPW_A = - 510,000 + 145,000\times \frac{1-(1+i)^{-11}}{i} + 10,000\times (1+i)^{-11}

No, we can determine the IRR using trial and error method. Assume rate = 25%

\large \implies NPW_A = - 510,000 + 145,000\times \frac{1-1.25^{-11}}{0.25} + 10,000\times 1.25^{-11}

On solving this we get

NPW = $ 21,037.37

Now assume rate = 27%

= NPWA = -510,000+145,000 x — 1-1.27-11 0.27 T10,000 X 1.27-11

Solve the above equation we get

NPW = - $ 10,982

Through linear interpolation we can determine the IRR

\large IRR_A = 25 + (27-25)\times \frac{NPW_{25\%}}{NPW_{25\%}- NPW_{27\%}}

\large IRR_A = 25 + 2\times \frac{21,037.37}{21,037.37- (-10,982)}

\large IRR_A = 25 + 2\times 0.657

\large IRR_A = 26.31\%

Solved it using excel got the same answer refer picture below

yr L A -510000 145000 145000 145000 145000 145000 145000 145000 145000 145000 145000 155000 26%

Now calculating IRR of alternative B

\large NPW_B= - 775,000 + 155,000(P/A,i\%,12) + 20,000(P/F,i\%,12)

Like we have calculated IRR for alternative A we can calculate for B

Assume rate = 15%

\large NPW_B= - 775,000 + 155,000(P/A,15\%,12) + 20,000(P/F,15\%,12)

On solving using the annuity table we get

NPW (15%) = $ 68,934.087

Now assume rate = 18%

\large NPW_B= - 775,000 + 155,000(P/A,18\%,12) + 20,000(P/F,18\%,12)

NPW(18%) = - $ 29,305.7562

Apply linear interpolation

\large IRR_B = 15 + (18 - 15)\times \frac{NPW_{15\%}}{NPW_{15\%}- NPW_{18\%}}

\large IRR_B = 15 + 3\times \frac{68,934.0878}{68,934.0878- (-29,305.7562)}

\large IRR_B = 15 + 3\times 0.70

\large IRR_B = 15 + 2.1

\large IRR_B = 17.1\%

Solved it using excel got the same answer refer attached picture below

IRR of Alternative A is higher than alternative B thus select A.

Select A.

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me please help. Thank you.